OpenStudy (anonymous):

Find a polynomial f(x) of degree 3 with real coefficients and the following zeros. 4, 2+i

OpenStudy (anonymous):

Well we know that the function has 2 zeroes and has is a third degree polynomial. So let's try making one by starting with the zeroes:\[y=(x-4)(x-(2+i))\]But in order for this to have real coefficients, we must in some way have to add another part that eliminates the imaginary parts from the expansion. We know for a fact that when complex numbers are multiplied by their conjugates, they result in real numbers. Can we put that knowledge to use here? @jmprz_793

OpenStudy (mertsj):

We know that complex answers come in pairs. If 2+i is an answer ten 2-i is also an answer.

OpenStudy (anonymous):

and so your job is to multiply \[(x-4)(x-(2+i))(x-(2-i))\] which is not as hard as it seems

OpenStudy (anonymous):

one way is just to multiply it out the other way is to work backwards to find \[(x-(2+i))(x-(2-i))\] you know the zero is \(2+i\) so you can put \[x=2+i\] subtract 2 \[x-2=i\] square \[(x-2)^2=-1\] expand \[x^2-4x+4=-1\] so \[x^2-4x+5=0\] is the original polynomial

OpenStudy (anonymous):

the last way is to memorize that if \(a+bi\) is a zero of a quadratic, the quadratic is \[x^2-2ax+(a^2+b^2)\]

OpenStudy (anonymous):

i am trying to enter the answer y=x^(2)-(6+i)x+(8+4i), but it does not let me enter the answer in parentheses @satellite73

OpenStudy (anonymous):

would it be y=x^(2)-6x-ix++4i

OpenStudy (mertsj):

The question is to find a polynomial f(x) Satellite told you that you job is to multiply : (x−4)(x−(2+i))(x−(2−i)) And then he proceeded to multiply the last two factors for you and told you that the product of (x-(2+i))(x-(2-i)) is x^2-4x+5 So finish the problem by multiplying (x-4)(x^2-4x+5)